Listed below are the annual tuition amounts of the 10 most expensive colleges in a country for a recent year. What does this​ "Top 10" list tell us about the population of all of that​ country's college​ tuitions?

$ 52 comma 247

  

$ 54 comma 223

  

$ 53 comma 376

  

$ 53 comma 666

  

$ 51 comma 123

$ 51 comma 215

  

$ 54 comma 121

  

$ 54 comma 121

  

$ 51 comma 990

  

$ 53 comma 063

Find the​ mean, midrange,​ median, and mode of the data set.
Answer #1

In order to solve this question I used R software.

R codes and output:

> x=c(52247,54223,53376,53666,51123,51215,54121,54121,51990,53063)
> summary(x)
Min. 1st Qu. Median Mean 3rd Qu. Max.
51120 52050 53220 52910 54010 54220
> table(x)
x
51123 51215 51990 52247 53063 53376 53666 54121 54223
1 1 1 1 1 1 1 2 1

Mean = 52910

Mid Range = (Max + min)/2 = (51120 + 54220) / 2 = 52670

Median = 53220

Mode = 54121